OpenStudy (amistre64):

does \[\frac{1}{1-2x}=2^n\]?

OpenStudy (anonymous):

not if x = 3!

OpenStudy (amistre64):

what are the steps to get it from the left to the right?

OpenStudy (amistre64):

apparently discrete math is my achilles heel lol

OpenStudy (anonymous):

this question honestly does not make any sense. you have an x on one side and n on the other.

OpenStudy (anonymous):

^^I agree.

OpenStudy (amistre64):

hmmm..... well, i think i write it right..

OpenStudy (amistre64):

should there be an 'n' up top instead of a 1?

OpenStudy (anonymous):

Then I think they are not equal.

OpenStudy (anonymous):

ooooooooooooohhhhhhhhh i bet i see what this should be. maybe. \[\frac{1}{1-x}=\Sigma x^n\] for -1<x<1 now if we replace x by 2x we get \[\frac{1}{1-2x}=\Sigma (2x)^n = \Sigma 2^nx^n\] for \[-\frac{1}{2}<x<\frac{1}{2}\] perhaps this is what you had in mind?

OpenStudy (amistre64):

that looks more like it :)